OpenStudy (anonymous):

A pyramid has a volume of 16 and an altitude of 5. What is the area of its base?

OpenStudy (ash2326):

\[\text{Volume of a Pyramid}=\frac{1}{3}\times \text{Area of Base}\times \text{height or altitude}\] Can you find area now?

OpenStudy (anonymous):

i need to find the area of the base and idk it

OpenStudy (ash2326):

We are given Volume=16 Altitude=5 Plug in these in this formula to find area \[\text{Volume of a Pyramid}=\frac{1}{3}\times \text{Area of Base}\times \text{height or altitude}\]

OpenStudy (ash2326):

Do you get this?

OpenStudy (anonymous):

yes i got it thanks

OpenStudy (ash2326):

good :D